SUPPOSE THAT DURING ANY HOUR IN A LARGE DEPARTMENT STORE, THE AVERAGE NUMBER OF SHOPPERS IS 448 , WITH A STANDARD DEVIATION OF 21 SHOPPERS. WHAT IS THE PROBABILITY OF RANDOMLY SELECTING 49 DIFFERENT SHOPPING HOURS, COUNTING THE SHOPPERS, AND HAVING THE SAMPLE MEAN FALL BETWEEN 441 AND 446 SHOPPERS, INCLUSIVE ?
Answer by stanbon(75887)

SUPPOSE THAT DURING ANY HOUR IN A LARGE DEPARTMENT STORE, THE AVERAGE NUMBER OF SHOPPERS IS 448 , WITH A STANDARD DEVIATION OF 21 SHOPPERS. WHAT IS THE PROBABILITY OF RANDOMLY SELECTING 49 DIFFERENT SHOPPING HOURS, COUNTING THE SHOPPERS, AND HAVING THE SAMPLE MEAN FALL BETWEEN 441 AND 446 SHOPPERS, INCLUSIVE ?
-----------------
z(441) = (441-448)/(21/sqrt(49)) = -7/7 = -1
z(446) = (446-448)/(21/sqrt(49)) = -2/7
=====
Ans:: P(441<= x <= 446) = P(-1<= z <=-2/7) = normalcdf(-1,-2/7) = 0.2289
-------------
Cheers,
Stan H.
------------